5. Tectonic Plating Inc. is considering a new three-year expansion project that re- quires an initial fixed asset investment of $1.65 million. The fixed asset falls into Class 10 for tax purposes (30%), and at the end of the three years can be sold for a salvage value of equal to its UCC. The project is estimated to generate $1,925,000 in annual sale, with costs of $595,000. If the tax rate is 40%, what is the OCF for each year of this project? Suppose the required return is 22%, what is the NPV? Furthermore, suppose the project requires an initial investment in net working capital of $245,000 and the fixed asset will have a market value of $450,000 at the end of the project. What is the new NPV?
Answer #1
Present Value (PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=Discount Rate=Required Return=22% 0.22
N=Year of Cash Flow
N YEAR 0 1 2 3
A Annual sale $1,925,000 $1,925,000 $1,925,000
B Costs $595,000 $595,000 $595,000
C Depreciation(30%) $495,000 $346,500 $242,550
D Accumulated depreciation $495,000 $841,500 $1,084,050
E=$1.65 m-B Undepreciated Capital Cost (UCC) $1,155,000 $808,500 $565,950
F=A-B-C Profit Before tax $835,000 $983,500 $1,087,450
G=F*0.4 Tax $334,000 $393,400 $434,980
H=F-G Profit after tax $501,000 $590,100 $652,470
I=H+C Operating Cash Flow (OCF) $996,000 $936,600 $895,020 SUM
K Present Value of Operating Cash Flow $816,393 $629,266 $492,894 $1,938,553
Present Value of Total Operating Cash flow $1,938,553
Initial cash flow -$1,650,000
Present Value of Salvage $311,673 (565950/(1.22^3)
Net Present Value (NPV) $600,226
CALCULATION OF NEW NPV
Cash Flow for net working capital ($245,000) $245,000
Salvage Value $450,000
Loss on salvage $115,950 (565950-450000)
Saving in Tax(40%) $46,380
Net Cash inflow on Salvage $496,380
Total additionalCash Flow in year3 $741,380 (245000+496380)
Present value of additional cash Inflow $408,283 (741380/(1.22^3)
Present Value of Total Operating Cash flow $1,938,553
Initial cash flow -$1,650,000
Additional working Capital investment ($245,000)
Present value of additional cash Inflow $408,283
Net Present Value (NPV) $451,837